# On Geometry and Topology of Hessian Manifolds

### v4

@@ 2.1 Obstructions from Sheaf Cohomology
Since the holonomy of $(M,\nabla)$ is trivial and the sheaf $\mathcal{P}$ of parallel 1-forms on $M$ is a locally constant sheaf of rank $n$, by abuse of notation we identify $\mathcal{P}$ with $\underline{\mathbb{R}}^{n}$. It is then readily seen that there exists a short exact sequence
@@
of cohomology groups. But since $M$ is simply connected, we have that the first de-Rham cohomology group of $M$ vanishes. Therefore, we obtain $\mathrm{H}^{1}(M;\mathcal{A})=0$. ∎
@@
Let $(M,\nabla)$ be an affine manifold, and let $\mathcal{A}$ be the sheaf of affine functions on $(M,\nabla)$. Suppose that $\pi\colon\tilde{M}\rightarrow M$ is a universal covering. Then $\pi^{-1}\mathcal{A}$ is the sheaf of affine functions on the affine manifold $(\tilde{M},\pi^{*}\nabla)$, and there exists a natural isomorphism
@@
Example 2.5.
There does exist Hessian manifold with finite but non-trivial fundamental group. The simplest example is the punctured affine space $\mathbb{R}^{3}-\{0\}$ quotient by the antipodal map $\textup{{x}}\mapsto-\textup{{x}}$, of which fundamental group is precisely $\mathbb{Z}/2\mathbb{Z}$.
@@
Let $(M,g)$ be a compact pseudo-Riemannian manifold, and let $\nabla$ be a connection on the cotangent bundle $T^{*}M\rightarrow M$. Then $g\neq\nabla(df)$ for all $f\in C^{\infty}(M)$.
@@
Let $(M,\nabla)$ be an affine manifold, and let $h$ be a Riemannian metric on $M$. The pair $((M,\nabla),h)$ is termed a Hessian manifold of Koszul type, if there exists a $1$-form $\eta\in\Omega^{1}(M)$ such that $h=\nabla\eta$, and in this case, the 1-form $\eta$ is termed a $1$-form potential for $((M,\nabla),h)$.
By Poincaré’s lemma, Hessian manifolds of Koszul type are indeed Hessian. As explained in the introduction section, this is, actually, the ultimate origin of the notion of Hessian structures.
2.2 Fundamental Group and Euler Number
@@
Moreover, for an affine manifold $(M,\nabla)$ with finite fundamental group, parallel 1-forms admit no periods as the first de-Rham cohomology group of $M$ vanishes. Therefore, the following notion of affine development is well-defined:
@@
In [14] and [13], Kostant et al. and Fried verified Conjecture 1.6 in some special cases.
Theorem 2.15 (Kostant-Sullivan,1975[13]).
Let $(M,\nabla)$ be a compact affine manifold, and let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. If $\nabla$ is geodesically complete, or equivalently the affine development $\Phi\colon\tilde{M}\rightarrow\mathbb{R}^{n}$ of $(\tilde{M},\pi^{*}\nabla)$ is a diffeomorphism, then the Euler characteristic $\chi(M)=0$.
Theorem 2.16 (Fried,1982[14]).
[…]
Combine Theorem 2.13 and Theorem 2.152.16, we immediately get the following result, confirming Conjecture 1.6 for Hessian manifolds.
Proposition 2.17.
The Euler characteristic of a compact Hessian manifold is zero.
Proof.
[…]
Corollary 2.7 shows that compact Hessian manifolds has infinite fundamental group. We shall now improve the result in Corollary 2.7.
@@
Let $\Omega$ be a convex domain in $\mathbb{R}^{d}$, and let $G$ be a group acting freely on $\Omega$. Then $G$ is torsion-free, that is, it contains no non-trivial element of finite order.
@@
$\textbf{b}:=\frac{1}{k}\sum_{i=0}^{k-1}g^{i}(\textbf{x})$
@@
The fundamental group a complete Hessian manifold is torsion-free.
@@
The fundamental group a complete Hessian manifold is either trivial or infinite.
@@
Let $((M,\nabla),h)$ be a compact Hessian manifold, and let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. Then $((\tilde{M},\pi^{*}\nabla),\pi^{*}h)$ is a complete Hessian manifold. Provided that the fundamental group of $M$ is finite, then $\tilde{M}$ is compact, which contradicts Theorem 2.13. This gives another proof of Corollary 2.7.
@@
By Proposition 2.19, the fundamental group $\pi_{1}(M)$ of $M$ is isomorphic to $\mathbb{Z}^{k}$ for some $k\in\mathbb{N}$. By Corollary 2.14 and the uniqueness of Eilenberg-MacLane space, $M$ is homotopy equivalent to the $k$-torus $\mathbb{T}^{k}$. Comparing the top homology groups of $M$ and $\mathbb{T}^{k}$ yields that $k=n$. The desired result then follows from Theorem A.3 in the appendix. ∎
@@ 2.3.1 Locally Constant Picard Group
Let $M$ be a differentiable manifold. The induced homomorphism $w_{1}\colon\mathrm{H}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}(M;\mathbb{Z}/2\mathbb{Z})$ in the long exact sequence
@@
Let $M$ be a differentiable manifold, and let $\varphi\colon\mathbb{R}^{\times}\rightarrow\mathbb{R}$ be the even extension of the logarithm function, that is, $\varphi(x)=\operatorname{log}|x|$ for all $x\in\mathbb{R}^{\times}$. The induced homomorphism $\operatorname{sh}:=\varphi_{*}\colon\mathrm{H}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}(M;\mathbb{R})$ in the long exact sequence
@@
$w_{1}\oplus\operatorname{sh}\colon\mathrm{H}^{1}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})\oplus\mathrm{H}^{1}(M;\mathbb{R}),\quad(\mathscr{L},D)\mapsto(w_{1}(\mathscr{L}),\operatorname{sh}(D))$
@@
In the case when the underlying manifold is aspherical, we may give an explicit description of the action of first Koszul class as an induced homomorphism.
Proposition 2.33.
Let $M$ be an aspherical differentiable manifold and $\pi_{1}(M)$ its fundamental group, so that we identify $\operatorname{Hom}(\pi_{1}(M),\mathbb{R})\equiv\mathrm{H}^{1}(M;\mathbb{R})$ via holonomy representation. Let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$, and let $\rho\colon\pi_{1}(M)\rightarrow\mathbb{R}^{\times}$ be its holonomy representation. Then, the homomorphism
$\pi_{1}(M)\rightarrow\mathbb{R},\quad g\mapsto\operatorname{log}|\rho(g)|$
as a group cohomology class in $\mathrm{H}^{1}(\pi_{1}(M);\mathbb{R})$ is precisely equal to the first Koszul class $\operatorname{sh}(D)$ of $(\mathscr{L},D)$.
Proof.
Denote by $G:=\pi_{1}(M)$ the fundamental group of $M$. Since $M$ is aspherical, the classifying map $M\to BG$ induces an isomorphism between the sheaf cohomology with locally constant coefficients and the corresponding group cohomology:
$\mathrm{H}^{1}(M;A)\cong\mathrm{H}^{1}(\pi_{1}(M);A)$
for any abelian group $A$.
Under this canonical isomorphism, the gauge equivalent class of $(\mathscr{L},D)$ in $\mathrm{H}^{1}(M;\mathbb{R}^{\times})$ corresponds to its holonomy representation $\rho\in\mathrm{H}^{1}(G;\mathbb{R}^{\times})\cong\operatorname{Hom}(G,\mathbb{R}^{\times})$.
Define group homomorphism $\varphi\colon\mathbb{R}^{\times}\rightarrow\mathbb{R}$ by $\varphi(x)=\operatorname{log}|x|$ for all $x\in\mathbb{R}^{\times}$. Then it readily seen that the following diagram commutes
$\begin{array}[]{ccc}\mathrm{H}^{1}(M;\mathbb{R}^{\times})&\overset{\operatorname{sh}}{\longrightarrow}&\mathrm{H}^{1}(M;\mathbb{R})\\ \downarrow&&\downarrow\\ \mathrm{H}^{1}(G;\mathbb{R}^{\times})&\overset{\varphi_{*}}{\longrightarrow}&\mathrm{H}^{1}(G;\mathbb{R})\end{array}$
and the desired property follows. ∎
We shall also be interested in a trivially criterion of elements of the locally constant Picard group. Recall that a line bundle is trivial if and only if it admits a nowhere vanishing global section. The following proposition gives a refinement in the case of flat $\mathbb{R}$-line bundles.
@@
To illustrate the power of the structure theorem of locally constant Picard group, we compare two proofs of the following torsion-free property.
@@
Let $s\in\mathrm{H}^{0}(M;\mathscr{L})$ be a nowhere-vanishing smooth section. Then by assumption and Proposition 2.34, there exists $f\in C^{\infty}(M)$ such that $\operatorname{exp}(kf)s^{\otimes k}$ is parallel. Then by Leibniz rule of differentiation,
$k(\operatorname{exp}(f)s)^{\otimes k-1}\otimes D(\operatorname{exp}(f)s)=D(\operatorname{exp}(kf)s^{\otimes k})=0$
and hence $D(\operatorname{exp}(f)s)=0$, as $k(\operatorname{exp}(f)s)^{\otimes k-1}$ is nowhere vanishing. Therefore $\operatorname{exp}(f)s$ is a nowhere vanishing parallel section of $(\mathscr{L},D)$. Again by Proposition 2.34, we conclude that $(\mathscr{L},D)$ is trivial as a flat $\mathbb{R}$-line bundle. ∎
Proof.
@@
2.3.2 Applications of Koszul Forms
@@
Recall that a Calabi-Yau manifold is a Kähler manifold with trivial canonical line bundle, or equivalently, admitting a holomorphic volume form. The following corollary exhibits an close analogy.
@@
Let $M$ be a differentiable manifold, and let $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ be a Hessian line bundle over $M$. Let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$, then we have $\operatorname{sh}(D)=[\varkappa]\in\mathrm{H}^{1}(M;\mathbb{R})$ is the first Koszul class of $(\mathscr{L},D)$.
@@
By the very definition, the first Koszul class $\operatorname{sh}(D)$ sends the class of the cocycle $\{g_{\alpha\beta}\}_{\alpha,\beta\in I}$ in $\mathrm{H}^{1}(M;\mathbb{R}^{\times})$ to the class of the cocycle $\{\operatorname{log}|g_{\alpha\beta}|\}_{\alpha,\beta\in I}$ in $\mathrm{H}^{1}(M;\mathbb{R})$.
@@
$\frac{1}{2}d\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
@@
$=\frac{1}{2}\operatorname{log}(\langle s_{\alpha}g_{\alpha\beta},s_{\alpha}g_{\alpha\beta}\rangle)-\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
@@
$=\frac{1}{2}\left(\operatorname{log}(g_{\alpha\beta}^{2})+\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)\right)-\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
$=\frac{1}{2}\operatorname{log}(g_{\alpha\beta}^{2})$
@@
Therefore, a Čech cocycle representing $\varkappa]$ is $\{\operatorname{log}|g_{\alpha\beta}|\}_{\alpha,\beta\in I}$, and hence $\operatorname{sh}(D)=[\varkappa]$. ∎
@@
for every vector field $X$ on $M$, where $|h|$ is the determinant of the matrix of which $(i,j)$-entry is $h(\partial/\partial x^{i},\partial/\partial x^{j})$ for some oriented affine coordinates $x^{1},\dots,x^{n}$ of the oriented affine manifold $(M,\nabla)$.

### v6

@@ 2.1 Methods of Sheaf Cohomology
Since the holonomy of $(M,\nabla)$ is trivial and the sheaf $\mathcal{P}$ of parallel 1-forms on $M$ is a locally constant sheaf of rank $n$, by abuse of notation we may identify $\mathcal{P}$ with $\underline{\mathbb{R}}^{n}$. It is then readily seen that there exists a short exact sequence
@@
of cohomology groups. But since $M$ is simply connected, we have that the first de Rham cohomology group of $M$ vanishes. Therefore, we obtain $\mathrm{H}^{1}(M;\mathcal{A})=0$. ∎
@@
Let $(M,\nabla)$ be an affine manifold, and let $\mathcal{A}$ be the sheaf of affine functions on $(M,\nabla)$. Suppose that $\pi\colon\tilde{M}\rightarrow M$ is the universal covering. Then $\pi^{-1}\mathcal{A}$ is the sheaf of affine functions on the affine manifold $(\tilde{M},\pi^{*}\nabla)$, and there exists a natural isomorphism
@@
Let $(M,g)$ be a compact pseudo-Riemannian manifold, and let $\nabla$ be a connection on the tangent bundle $TM\rightarrow M$. Then $g\neq\nabla(df)$ for all $f\in C^{\infty}(M)$.
@@
Let $(M,\nabla)$ be an affine manifold, and $h$ a Riemannian metric on $M$. The pair $((M,\nabla),h)$ is termed a Hessian manifold of Koszul type, if there exists a $1$-form $\eta\in\Omega^{1}(M)$ such that $h=\nabla\eta$, and in this case, the 1-form $\eta$ is termed a $1$-form potential for $((M,\nabla),h)$.
By Poincaré lemma, Hessian manifolds of Koszul type are indeed Hessian. As explained in the introduction section, this is, actually, the ultimate origin of the notion of Hessian structures.
2.2 More on Fundamental Group
@@
Moreover, for an affine manifold $(M,\nabla)$ with finite fundamental group, parallel 1-forms admit no periods as the first de Rham cohomology group of $M$ vanishes. Therefore, the following notion of affine development is well-defined:
@@
Corollary 2.6 shows that compact Hessian manifolds have infinite fundamental group. We shall now improve the result in Corollary 2.6.
@@
Let $\Omega$ be a convex domain in $\mathbb{R}^{d}$, and $G$ a group acting freely on $\Omega$. Then $G$ is torsion-free, that is, it contains no non-trivial element of finite order.
@@
$\textbf{b}:=\frac{1}{k}(g(\textbf{x})+g^{2}(\textbf{x})+\cdots+g^{k}(\textbf{x}))$
@@
The fundamental group of a complete Hessian manifold is torsion-free.
@@
The fundamental group of a complete Hessian manifold is either trivial or infinite.
@@
Let $((M,\nabla),h)$ be a compact Hessian manifold, and let $\pi\colon\tilde{M}\rightarrow M$ be the universal covering of $M$. Then $((\tilde{M},\pi^{*}\nabla),\pi^{*}h)$ is a complete Hessian manifold. Provided that the fundamental group of $M$ is finite, then $\tilde{M}$ is compact, which contradicts Theorem 2.12. This gives another proof of Corollary 2.6.
@@
By Proposition 2.16, the fundamental group $\pi_{1}(M)$ of $M$ is isomorphic to $\mathbb{Z}^{k}$ for some $k\in\mathbb{N}$. By Corollary 2.14 and the uniqueness of Eilenberg-MacLane space, $M$ is homotopy equivalent to the $k$-torus $\mathbb{T}^{k}$. Comparing the top homology groups of $M$ and $\mathbb{T}^{k}$ yields that $k=n$. The desired result then follows from the Borel conjecture, which is proved for homotopy tori. See for example [15] and [8]. ∎
@@ 2.3 Flat Line Bundles
Let $M$ be a differentiable manifold. The induced homomorphism
$w_{1}\colon\mathrm{H}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}(M;\mathbb{Z}/2\mathbb{Z})$
in the long exact sequence
@@
Let $M$ be a differentiable manifold, and let $\varphi\colon\mathbb{R}^{\times}\rightarrow\mathbb{R}$ be the even extension of the logarithm function, that is, $\varphi(x)=\operatorname{log}|x|$ for all $x\in\mathbb{R}^{\times}$. The induced homomorphism $\kappa_{1}:=\varphi_{*}\colon\mathrm{H}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}(M;\mathbb{R})$ in the long exact sequence
@@
$w_{1}\oplus\kappa_{1}\colon\mathrm{H}^{1}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})\oplus\mathrm{H}^{1}(M;\mathbb{R}),\quad(\mathscr{L},D)\mapsto(w_{1}(\mathscr{L}),\kappa_{1}(D))$
@@
We shall also be interested in a triviality criterion of elements of the locally constant Picard group. Recall that a line bundle is trivial if and only if it admits a nowhere vanishing global section. The following proposition gives a refinement in the case of flat $\mathbb{R}$-line bundles.
@@
To illustrate the power of the structure theorem of locally constant Picard group, we present a proof of the following torsion-free property.
@@
2.4 Applications of Koszul Forms
@@
Recall that a Calabi-Yau manifold is a Kähler manifold with trivial canonical line bundle, or equivalently, admitting a holomorphic volume form. The following corollary exhibits a close analogy.
@@
Let $M$ be a differentiable manifold, and let $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ be a Hessian line bundle over $M$. Let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$, then we have $\kappa_{1}(D)=[\varkappa]\in\mathrm{H}^{1}(M;\mathbb{R})$ is the first Koszul class of $(\mathscr{L},D)$.
@@
By the very definition, the first Koszul class $\kappa_{1}(D)$ sends the class of the cocycle $\{g_{\alpha\beta}\}_{\alpha,\beta\in I}$ in $\mathrm{H}^{1}(M;\mathbb{R}^{\times})$ to the class of the cocycle $\{\operatorname{log}|g_{\alpha\beta}|\}_{\alpha,\beta\in I}$ in $\mathrm{H}^{1}(M;\mathbb{R})$.
@@
$\varkappa_{\alpha}:=\frac{1}{2}d\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
@@
Therefore, a Čech cocycle representing $\varkappa]$ is $\{\operatorname{log}|g_{\alpha\beta}|\}_{\alpha,\beta\in I}$, and hence $\kappa_{1}(D)=[\varkappa]$. ∎
@@
for every vector field $X$ on $M$, where $|h|$ is the determinant of the matrix of which $(i,j)$-entry is $h(\partial/\partial x^{i},\partial/\partial x^{j})$ for some affine coordinates $x^{1},\dots,x^{n}$ of the oriented affine manifold $(M,\nabla)$.